RC Meeting and Run Plan, Thursday, January 9th 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Target Readiness

  • Polarimeter acquired and DOC measurements made. The angles were corrected and longitudinal and transverse have very high DOC polarization!
  • Spin up measurements conducted this morning and they indicate towards growing polarization (but limited by data points). What tests should we perform before closing up?
  • Collimators are being worked in in the machine shop. Expected to be done today evening.
  • Stepper motor had homing issues and would go to random angles. The problem disappeared after rebooting.

Remaining work

In the enclosure:

  • Installing NMR legs and 2 BNC cables.
  • Install EPR fiber bundle cable
  • Install Reference cell RTDs
  • Check the long. + trans. spectral analyses + mounts
  • Cooling jets alignments
  • Reference cell gas connections
  • Target ladder alignment and adjusting the software encoder positions (if needed)
  • Install collimators
  • Follow the enclosure backout checklist
  • Check the angles on QWP in the laser enclosure.

Outside the enclosure:

  • Interlock wire broken - need tech's help
  • General clean up around the pivot area
  • Follow the hall backout checklist

Current Run Plan

  • Thursday: Continue with spin up tests
  • Friday: Continue with spin up tests + follow remaining tasks listed above. Close up the target enclosure.
  • Saturday: Start taking 1-pass data for two days with the same settings as in December.
  • Sunday: 1-pass data
  • Monday: Remove collimators and close the panels. Make a pass change.

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[1]
  • Temporary Laser instructions can be found here [2]

Ongoing Work in the Hall

Planned Accelerator Configuration Changes

Special Requests

  • Keep target expert schedule filled out on whiteboard